### v1

Abstract
Agentic systems offer a potential path to solve complex clinical tasks through collaboration among specialized agents, augmented by tool use and external knowledge bases. Nevertheless, for chest X-ray (CXR) interpretation, prevailing methods remain limited: (i) reasoning is frequently neither clinically interpretable nor aligned with guidelines, reflecting mere aggregation of tool outputs; (ii) multimodal evidence is insufficiently fused, yielding text-only rationales that are not visually grounded; and (iii) systems rarely detect or resolve cross-tool inconsistencies and provide no principled verification mechanisms. To bridge the above gaps, we present RadAgents, a multi-agent framework for CXR interpretation that couples clinical priors with task-aware multimodal reasoning. In addition, we integrate grounding and multimodal retrieval-augmentation to verify and resolve context conflicts, resulting in outputs that are more reliable, transparent, and consistent with clinical practice.
keywords:
Multi-agent system, multimodal reasoning, chest X-ray, image interpretation.
Data and Code Availability
We use the following public datasets: MIMIC-CXR-JPG (Johnson et al., 2019), MS-CXR (Boecking et al., 2022), and MS-CXR-T (Bannur et al., 2023), which are accessible under their respective data use agreements. We plan to release the code after obtaining organizational approval.
Institutional Review Board (IRB)
This work does not require IRB approval.
1 Introduction
Figure 1: RadAgents framework. Each ABCDE subagent executes in parallel guided by clinical workflows, lowering latency, preserving isolation to avoid long-context drift, and improving trustworthiness.
Chest X-ray (CXR) imaging is a cornerstone of pulmonary screening, diagnosis, and follow-up, accounting for the largest share of diagnostic radiology examinations performed worldwide each year (Cid et al., 2024). Yet systematic assessment of thoracic structures remains labor-intensive, imposing a substantial time burden on radiologists (Fallahpour et al., 2025). The gradual introduction of AI into clinical practice shows promise for alleviating this workload (Zhang et al., 2024; Tanno et al., 2025). However, prevailing systems fall short on complex multimodal reasoning, such as integrating findings across disparate image regions, which is central to radiologists’ practice; most adhere to end-to-end designs in which the visual encoder executes a single, front-end pass and subsequent reasoning proceeds purely in text (Wang et al., 2025). This encode-once, text-only paradigm decouples the reasoning trajectory from evolving visual evidence, leading to failures on tasks that require iterative re-inspection, precise measurements, and cross-comparisons (Liu et al., 2025).
A promising path for clinical reasoning is to augment large language models, including multimodal variants, with external tools. By delegating perceptual and classification subtasks such as organ or region segmentation and disease classification to validated modules, the language model can focus on planning and synthesis. Several agentic frameworks already explore this idea, from fine tuned tool use (Li et al., 2024; Nath et al., 2025) to pipelines that call commercial models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), although multiple agent coordination and/or debate introduces considerable computational overhead. In CXR interpretation, RadFabric (Chen et al., 2025) integrates diagnostic agents with a separate reasoning agent, and MedRAX (Fallahpour et al., 2025) expands task coverage by incorporating additional task specific models. Despite gains over single-model baselines, integration and reasoning steps are often opaque and not aligned with clinical workflow, which undermines trust and creates safety risks, and most systems still lack explicit verification and conflict resolution.
In this paper, we introduce RadAgents, a training free framework (that is, no weight updates), designed to be trustworthy and extensible for CXR interpretation, especially in cases that require complex multimodal reasoning. Evaluated on three datasets across two core tasks: report generation and visual question answering (VQA), RadAgents achieves state of the art performance, surpassing the strongest baselines by 10.2% on MS-CXR, 29.6% on MIMIC-CXR, and 21.5% on MS-CXR-T, underscoring the value of incorporating clinically-aware workflows. Our agentic design, enhanced with multimodal retrieval-augmentation, delivers consistent gains: an average improvement of 8.0% over the no-retrieval variant, demonstrating its effectiveness in mitigating context conflicts and producing more reliable results.
2 Methods
RadAgents is a multi agent system with seven specialized agents (Figure 1). Five implement the clinical ABCDE review scheme (Hodler et al., 2019): Airway, Breathing, Circulation, Diaphragm, and Everything else. In addition, an Orchestrator agent analyzes each query and routes tasks to the appropriate specialists with the required patient context (for example, imaging view and prior studies), and a Synthesizer agent integrates their outputs, resolves conflicts, and produces the final output.
This design confines context to task specific compartments, reducing the information each agent must process and simplifying context compression for synthesis. It also allows parallel execution, lowering latency for long reasoning. For common and clinically significant CXR findings such as cardiomegaly and pleural effusion, we curate radiologist like workflows that guide tool selection and clinically grounded reasoning (see the demonstration example in Appendix A). Beyond these predefined templates, the language model handles queries that fall outside fixed workflows, plans tool use, and enables flexibility in interpretation.
2.1 Task-aware Sub-Agents
Each sub-agent, also called the ABCDE agent, has a defined purpose and domain of expertise. Each is governed by a custom system prompt derived from clinical guidelines and maintains its own context window. The main scope and objectives of them are:
Airway agent: Systematically assess the central thorax for airway patency, alignment, and paratracheal lesions; for example, determine tracheal position (midline versus deviation).
Breathing agent: Survey the lungs and pleura for parenchymal and pleural pathology; for example, detect opacities (alveolar, interstitial, nodular), infiltrates, and consolidation patterns.
Circulation agent: Evaluate the cardiac silhouette, mediastinum, and vessels; for example, compute the cardiothoracic ratio.
Diaphragm agent: Assess diaphragmatic integrity and look for subdiaphragmatic air; for example, compare right and left diaphragm height.
Everything Else agent: Identify other chest wall, soft tissue, and foreign body findings.
2.2 Radiologist-like Workflows
Solving different radiological tasks requires distinct reasoning modes and corresponding evidence. Figure 6 shows that multimodal reasoning, particularly interleaved quantitative measurements, strengthens reasoning for enlarged heart detection, whereas a simple grounding guided approach fails. Motivated by this, RadAgents curates clinically driven subtasks and workflow templates within the subagents’ system prompts, organized into five modes: (M1) measurement, (M2) localization, (M3) characterization, (M4) relational and comparative reasoning, and (M5) diagnosis. Complex clinical reasoning composes these modes and triggers the appropriate tool calls within the agentic system; for example, judging progression of effusion volume uses M1 and M4. A detailed description of workflows can be found in Appendix E.
2.3 Global Controller Module
The global controller comprises the Orchestrator and the Synthesizer. The Orchestrator selects subagents and allocates tasks with appropriate patient context, and the Synthesizer integrates their outputs, verifies consistency, and resolves errors and conflicts. The major components are detailed below.
Query analysis. Given a query, the Orchestrator drafts a high level plan, selects the relevant subagents, and chooses the reasoning mode: ReAct when no workflow is specified (Yao et al., 2023), or Plan-and-Execute (P&E) (WANG et al., 2023) when a workflow template is available. This keeps the system language driven and adaptable across queries.
Tool cards. We employ a suite of models as tools for distinct CXR tasks: CheXagent (Chen et al., 2024b) for VQA, MAIRA-2 (Bannur et al., 2024) for grounding, the CheXpert Plus report generator (Chambon et al., 2024), and classification and organ segmentation models from TorchXRayVision (Cohen et al., 2022). In addition, we include unique programming tools that return zoomed-in quarter patches or serve for measurement and calculation purposes.
Context verifier for tools. No tool is perfect, as their capabilities are constrained by model size and training data. When uncertainty arises, we trigger a verification step in which an advanced multimodal LLM serves as a judge (Chen et al., 2024a), filtering out incorrect outputs such as erroneous masks.
Retrieval-augmented conflic resolution. Tool outputs can conflict. On the Synthesizer side, we apply Visual Retrieval-Augmented Generation (V-RAG) (Chu et al., 2025): the agent retrieves clinically similar chest radiographs (based on image embeddings from Rad-DINO (Perez-Garcia et al., 2025)) and accompanying context like patient notes and uses these exemplars to adjudicate discrepancies among tools (Figure 2). This mirrors routine radiologic practice, in which clinicians consult similar cases and content to calibrate interpretation.
Figure 2: Resolving the conflicts via V-RAG.
3 Experiments
3.1 Experimental Setup
To demonstrate the generality of RadAgents, we evaluate it on three tasks with increasing reasoning complexity: VQA for existence and attributes (E&A), VQA for comparison and progression (C&P), and report generation. The data statistic and details are shown in Table 2.
Baselines. We instantiate all agents with GPT-4o. For comparison, we include (1) GPT-4o (Hurst et al., 2024), (2) GPT-4o with ReAct, where tools are available via function calling but without explicit workflow steering, and (3) two medical specialist models, CheXagent and MedGemma (Sellergren et al., 2025). Unless otherwise noted, the number of retrieved exemplars for V-RAG is set to $k=3$ (see Appendix F for an ablation on $k$). We report results for two variants of RadAgents, with and without V-RAG.
Metrics. For VQA (E&A) and report generation, we use standard CXR text metrics (explanation in Appendix B): RadGraph F1 (Jain et al., 2021), CheXbert macro F1 across 14 labels (Smit et al., 2020), RaTE (Zhao et al., 2024), and GREEN (Ostmeier et al., 2024). Because the outputs are sentences, these metrics capture clinical correctness and consistency. For VQA (C&P), where the output is one of three choices, we report accuracy.
Table 1: Evaluation on VQA (E&A) and report generation with baselines and RadAgents.
Method CheXbert-macro-F1(14) RadGraph-F1 RaTE GREEN Avg.
[…]
CheXagent $0.3321$ $0.1817$ $0.4526$ $0.3429$ $0.3273$
MedGemma $0.3827$ $0.1624$ $0.5648$ $0.5723$ $0.4205$
GPT-4o $0.3219$ $0.0928$ $0.4122$ $0.2127$ $0.2599$
GPT-4o w/ ReAct $0.3613$ $0.1221$ $0.5034$ $0.4619$ $0.3622$
RadAgents wo/ V-RAG $0.4128$ $0.1925$ $0.5147$ $0.5841$ $0.4260$
RadAgents $0.4426$ $0.2314$ $0.5758$ $0.6032$ 0.4632
[…]
CheXagent $0.2916$ $0.1318$ $0.4129$ $0.1825$ $0.2547$
MedGemma $0.2413$ $0.1189$ $0.4728$ $0.2416$ $0.2686$
GPT-4o $0.2237$ $0.1324$ $0.4635$ $0.3138$ $0.2833$
GPT-4o w/ ReAct $0.3521$ $0.1329$ $0.4731$ $0.3325$ $0.3226$
RadAgents wo/ V-RAG $0.4412$ $0.1826$ $0.5238$ $0.3821$ $0.3824$
RadAgents $0.4727$ $0.1829$ $0.5644$ $0.4527$ 0.4182
3.2 Existence and Attributes
The VQA questions cover seven common findings in CXR: atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, and pneumothorax, derived from the standard test split of MS–CXR. Each image receives the following prompt: {infobox}<image> Describe if [finding] is present; if present, describe [attributes].
Table 3.1 shows that adding V-RAG to the agent improves all metrics over the ablation: +0.0298 CheXbert, +0.0389 RadGraph, +0.0611 RaTE, +0.0191 GREEN, raising Avg. from 0.4260 to 0.4632 (+0.0372, +8.7%). Among baselines, MedGemma is strongest (0.4205 Avg.) yet remains 10.2% below RadAgents. GPT-4o benefits from ReAct (+0.1023 Avg.) but still trails the full agent by 0.1010.
3.3 Comparison and Progression
We use MS–CXR–T to assess stability, improvement, or worsening of a specific positive finding (consolidation, edema, pleural effusion, or pneumothorax). We only retain cases where the metadata indicates a consensus among human reviewers. We pose a comparative question that explicitly references the prior study. The prompt template is: {infobox}Given current image <image>, and previous image <image>, decide if [finding] is improving, stable, or worsening.
Figure 3 shows that RadAgents+V-RAG achieves the best overall accuracy on MS-CXR-T, surpassing the ablated agent and all LLM baselines. The ordering mirrors E&A, underscoring the value of retrieval over similar studies for longitudinal reasoning.
3.4 Report Generation
We construct a MIMIC–CXR subset aligned with MS–CXR identities so that findings queried in VQA are represented in the corresponding reports. Prompts request generation of the Findings section, and all agents are activated by default. The prompt combines the template from Section 3.2 with a curated list of clinically significant findings, following prior work (Tu et al., 2024; Peng et al., 2025); details appear in Appendix D.
Table 3.1 shows that, RadAgents attains the best report quality (Avg. 0.4182). V-RAG contributes +0.0335 Avg. (+8.7%), with the largest gain on GREEN (+0.0706), plus lifts on CheXbert (+0.0315) and RaTE (+0.0406). GPT-4o+ReAct improves over GPT-4o (+0.0393) but remains 0.0956 below RadAgents, while MedGemma drops on this task (0.2686 Avg.).
Figure 3: VQA results regrading progression.
4 Discussion
In this work, we propose the first radiologist-like agentic system, which demonstrates superior performance in CXR interpretation. Further improvements are required for broader application: (1) the context verifier can only detect errors but cannot modify outputs, especially for visual evidence such as bounding boxes, so stronger supporting models are needed; (2) current experiments are limited to frontal-view images, as most existing tools are incompatible with lateral views; (3) prompting strategies could be further optimized; and (4) the framework could be extended to additional modalities, such as CT and MRI.
Acknowledgments
We would like to thank other members of Oracle Health AI for their support while developing our system, and Raefer Gabriel, Sri Gadde, Mark Johnson, Devashish Khatwani, Yuan-Fang Li, Anit Sahu, Praphul Singh, and Vishal Vishnoi for insightful feedback and discussions.
Appendix A Demonstrations
A.1 Full reasoning trajectory for tracheal deviation detection
Figure 5 shows the reasoning trajectory for the query “Is the trachea midline?” The sequence proceeds as follows: (a) the orchestrator agent analyzes the query, activates the airway agent, and initiates actions; (b) the returned tool context is visualized; (c) the synthesizer agent integrates the context but issues a warning due to low-confidence segmentation results, triggering the context verifier—specifically, GPT-4o is invoked to re-judge and answer the query; and (d) synthesis is completed, producing the final answer with an associated confidence score.
A.2 Failure case in CTR calculation
Most measurement reasoning relies on segmentation or grounding tools for support. For example, measuring heart width directly involves counting the number of pixels between boundary points. However, due to the high cost of annotation, ground-truth datasets for training robust segmentation or grounding models are limited. Moreover, most organ segmentation datasets are curated from normal images; when abnormalities obscure the organs, segmentation often fails, leading to downstream measurement errors. Figure 4(b) illustrates the segmentation masks for a normal CXR and a CXR with effusion. In the effusion case, both heart width and thoracic width are measured incorrectly due to inappropriate masks. These results further emphasize the need for a context verifier and resolver.
Figure 4: CTR calculation with segmentation masks on (a) normal and (b) effusion cases. Inaccurate masks in the abnormal case lead to incorrect heart and thoracic width measurements.
Figure 5: RadAgents’ reasoning trajectory for tracheal deviation detection.
Appendix B Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (Smit et al., 2020), RadGraph (Jain et al., 2021), GREEN (Ostmeier et al., 2024), and RaTEScore (Zhao et al., 2024), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
Appendix C Dataset
Operating an agentic system is costly because multimodal multi-step reasoning entails many LLM API calls; accordingly, we did not conduct very large scale experiments. We evaluate on MS–CXR and MS–CXR–T, whose annotations enable verification of intermediate outputs and support future reinforcement learning to incentivize tool use with open source models. For report generation, we condition on the current study and a single prior frontal image, rather than the full imaging history, to limit context length and processing cost. Dataset statistics are shown in Table 6.
Dataset # Cases # Images per case Has prior?
MIMIC-CXR (subset) 181 2 Yes
MS-CXR (test set) 181 1 No
MS-CXR-T 785 2 Yes
Table 2: Details of datasets used in RadAgents.
Appendix D Prompting
The input to RadAgents includes not only the image and query but also optional clinical context, such as patient demographics, indication, acquisition technique, comparison studies, and examination details (e.g., view/projection and region). We denote this aggregated context as Context in the templates for comprehensive CXR scanning and for report generation.
{infobox}
Instruction: You are a helpful radiology assistant. Describe what lines, tubes, and devices are present and each of their locations. Describe if pneumothorax is present; if present, describe size on each side. Describe if pleural eusion is present; if present, describe amount on each side. Describe if lung opacity (atelectasis, brosis, consolidation, inltrate, lung mass, pneumonia, pulmonary edema) is present; if present, describe kinds and locations. Describe the cardiac silhouee size. Describe the width and contours of the mediastinum. Describe if hilar enlargement is present; if enlarged, describe side. Describe what fractures or other skeletal abnormalities are present.
Given the [view] X-ray image(s) <images>,
Q: Describe the findings in the image following the Instructions, and Context: {context}.
Figure 6: Different queries should trigger different reasoning modes. Simply cropping regions of interest and curating visual chain-of-thought reasoning is not a panacea.
Appendix E Reasoning modes and agentic workflows
We first introduce clinically aware interleaved reasoning, which identifies significant findings through context assembly from metadata, targeted tool use, and self guided reasoning. We then outline the RadAgents workflow, driven by prompt based task decomposition, strategic tool selection, and report synthesis.
Why interleaved reasoning?
Classical vision and language systems compress the image once and then reason only in text. This single pass approach is ill suited to radiology, where clinicians iteratively survey the study, escalate suspicious findings to measurement, revisit earlier impressions as new evidence appears, and compare with priors. RadAgents operationalizes this practice by interleaving perception and reasoning: at intermediate points the system inspects additional visual evidence (for example, cropped regions or segmentation overlays) and updates its hypothesis before proceeding. A simple protocol is a two phase visual chain of thought (VCoT): first, a pure visual assessment that answers the question and cites observable evidence without tool outputs; second, an evidence validation step that reveals tool conclusions (for example, a measurement or a mask) and records agreement, disagreement, or uncertainty with a recalibrated confidence. This improves transparency and reduces anchoring on imperfect tools. Figure 6 illustrates how interleaved quantitative evidence aids enlarged heart detection, whereas a simple grounding guided approach fails.
(M1) Measurement
Goal. Provide objective and reproducible judgments for geometry constrained findings.
When. Explicit measurement requests or size abnormalities suggested by a sweep.
Tools. Segmentation yields organ masks and derived metrics; the cardiothoracic ratio (CTR) uses maximal cardiac width and thoracic width from bilateral lung extents, with projection recorded (PA versus AP) and raw pixel widths logged.
Evidence. Numeric values with projection, overlay visuals of the masks used, and brief caveats.
(M2) Localization
Goal. Localize small or subtle targets that benefit from high resolution crops.
When. Explicit localization queries or equivocal global signals.
Tools. Grounding proposes bounding boxes; segmentation constrains search when organ context matters; crop policies adapt to target scale.
Evidence. Boxes with confidence, region thumbnails, and landmark distances when relevant.
(M3) Characterization
Goal. Describe texture, morphology, distribution, and severity of parenchymal and pleural findings.
When. Opacity related queries or when a sweep suggests edema, atelectasis, pneumonia, or fibrosis.
Tools. Zone prioritization by a classifier, lung masks to focus attention, region crops from suspicious zones, VQA to standardize descriptors, and VCoT to justify labels.
Evidence. Pattern and distribution labels with severity, plus representative crops tied to the cited features.
(M4) Relational and comparative reasoning
Goal. Determine improvement, worsening, or stability relative to a prior study, as well as the relationships between each finding like complication.
When. Explicit comparison or association requests or any mention of a prior.
Tools. Temporal alignment of regions, adjustment for projection or view differences when possible, signed change statements in VCoT, and measurement deltas for quantitative targets.
Evidence. Paired crops and a concise comparison statement.
(M5) Differential Diagnosis
Goal. Synthesize evidence across modes to issue a patient level conclusion and differential with calibrated confidence.
When. After upstream modes have supplied sufficient evidence or when the query explicitly seeks a diagnosis.
Tools. The synthesizer aggregates structured findings, checks consistency, reconciles conflicts, and maps evidence to diagnostic statements aligned with clinical guidance.
Evidence. A short justification that links key measurements, locations, and patterns to the final conclusion, plus uncertainty notes when appropriate.
Appendix F Details of V-RAG
Multimodal retrieval.
We retrieve images and their textual descriptions that align with the features of target medical images following (Chu et al., 2025). These references, rich in visual and textual medical details, guide response generation. To obtain embeddings, we use Rad-DINO, which provides robust representations across diverse CXR image types. For each image $X_{img}$, we extract its embedding $E_{img}=\mathbf{R}^{d}$, with $d=768$, and store them in the embedding memory $\mathcal{M}$.
For efficient retrieval during inference, we build $\mathcal{M}$ using FAISS  [^5] https://github.com/facebookresearch/faiss, a GPU-accelerated vector search system. We employ approximate kNN with the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2018), enabling retrieval of the top-$k$ most similar images in $\mathcal{M}$.
Table 3: Retrieval quality at different $k$ values, showing the trade-off between helpful and harmful rates.
$k$ Helpful-rate Harmful-rate
1 0.48 0.09
3 0.62 0.14
5 0.65 0.20
Sensitivity of $k$.
We study how retrieval quality changes with the number of retrieved studies $k$ with the sampled 100 cases from the MS-CXR test set used in this study. For each setting, we compute two metrics: helpful-rate, the percentage of retrieved studies that improve the answer, and harmful-rate, the percentage that hurt the answer. As shown in Table 3, the helpful-rate increases from 0.48 at $k=1$ to 0.65 at $k=5$, while the harmful-rate also rises from 0.09 to 0.20. This illustrates a trade-off: retrieving more studies provides greater chances of including helpful evidence but also increases the risk of introducing misleading content. To balance these effects, we choose $k=3$ by default, achieving a helpful-rate of 0.62 with a moderate harmful-rate of 0.14.
Augmented Inference.
In the inference stage, we encode the query image $X_{q}$ to obtain its embedding. We then retrieve the top-$k$ most similar images from $\mathcal{M}$, represented as $(I_{1},\dots,I_{k})$ with their corresponding reports $(R_{1},\dots,R_{k})$. These references are appended to the input of multimodal LLM to guide generation. The prompt is structured as:
{infobox}
This is the $i$-th similar image and its report for your reference. [Reference]i ... According to the query image and the references, [Question] [Query Image]. where each reference is denoted as $(I_{i},R_{i})$.